paopao03:

Part A: The area of a square is (16x2 + 24x + 9) square units. Determine the length of each side of the square by factoring the area expression completely. Show your work. (5 points) Part B: The area of a rectangle is (4x2 − 49y2) square units. Determine the dimensions of the rectangle by factoring the area expression completely. Show your work. (5 points) (10 points)

mdobbs6856:

Part A 9x^2 + 24x + 16 = (3x + 4)(3x + 4) length of each side =? Par B 16x^2 - 25y^2 = (4x + 5y )( 4x - 5y) length =?

supie:

Yeah, I think that is right For part A . . . \[9x^2 + 24x + 16= (3x + 4)(3x + 4)\] So with that you can find the length of all sides. And for part B \[16x^2 - 25y^2= (4x + 5y )( 4x - 5y)\] So now we can find the length and width \[length=?\\width=?\]